OpenStudy (anonymous):

Find each value. Round to the nearest tenth if necessary. 8.2 is 2% of what number

OpenStudy (anonymous):

ok so here's how to do these problems

OpenStudy (anonymous):

i think its 4.1 but im not quite sure

OpenStudy (anonymous):

say number is X...now 2% of x is 8.2 i.e. X*0.02=8.2 so x= 8.2/0.02

OpenStudy (anonymous):

it should be 410 i guess

OpenStudy (anonymous):

o ok i got it now thanks

OpenStudy (anonymous):

how would you do this problem: Find each value. Round to the nearest tenth if necessary. What percent of 9 is 27?......would you multiply 27 by 9

OpenStudy (anonymous):

ok let's say the answer is X...so x % of 9 is 27 (X/100)*9=27 x/100=27/9 x/100=3 x=300

OpenStudy (anonymous):

o so you would do 27 divided by 9 then multiply by 100

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

what about this 1 : Find each value. Round to the nearest tenth if necessary. 210% of 30 would you divide and multiply by 100

OpenStudy (anonymous):

ok here's the thing...whenever you are asked find percentage of something, replace the % sign with 1/100 and the multiply with the number whose % u want

OpenStudy (anonymous):

did u get it?

OpenStudy (anonymous):

so you would replace 210 with

OpenStudy (anonymous):

thats the part where I dont get what would i put there

OpenStudy (anonymous):

u don't replace the number!...replace the % sign...so here u r asked 210 % of 30...so u replace % by 1/100...u get 210/100 and u multiply this no with 30

OpenStudy (anonymous):

now u get it?

OpenStudy (anonymous):

yeah and i didn't mean replace 210 i just typed it cause i wasn't thinking i meant with the percent

OpenStudy (anonymous):